12.0 m )j . What are (a) the Here are two vectors: а-(6.00 m )i-4.50 m lj and b-| 9.00 m li + magnitude and (b) the angle (counterclockwise from the axis defined by ) of a ? what are (c) the magnitude and (d) the angle of b ? What are (e) the magnitude and (f) the angle of a + b ; (g) the magnitude and (h) the angle of b-a and (i) the magnitude and () the angle of a -b (State your angle as a positive number.) (k) What is the angle between the directions of b -a and a-b Units (a) Number (b) Number (c) Number (d) Number (e) Number (f) Number (9) Number (h) Number (i) Number Units Units Units Units Units Units Units Units 6i) Number Units
Answer #1

given

a = 6.00 i - 4.50 j

b = 9.00 i + 12.0 j

a )

| a | = ( 62 + (-4.5)2 )1/2

= 7.5 m

b )

angle = tan-1 ( -4.5/6 )

= - 36.87o

here 360 - 36.87 = 323.13o anti clock wise from the x - axis

c )

| b | = ( 92 + 122 )1/2

= 15 m

d )

angle = tan-1 ( 12/9 )

= 53.13o

here 53.13o anti clock wise from the x - axis

e )

a + b = ( 6 + 9 ) i + ( -4.5 + 12 ) j

= 15 i + 7.5 j

| a + b | = ( 152 + 7.52 )1/2

= 16.77 m

f )

angle is = tan-1 ( 7.5/15 )

= 26.56o

here 26.56o anti clock wise from the x - axis

g )

| b - a | = | ( 6 - 9 ) i | + | ( 12 + 4.5 )j |

= 3 i + 16.5 j

| b - a | = ( 32 + 16.52 )1/2

= 16.77 m

h )

angle = tan-1 ( 16.5/3 )

= 79.695o anti clock wise from x - axis

i )

a - b = - 3 i - 16.5 j

| a - b | = 16.77 m

j )

angle = tan-1 ( - 16.5 / -3 )

= 79.695o

it is now 360 - 79.695 = 259.695o

k )

so the angle is = 259.695 - 79.695

= 180o